— HOV NAI 2K 385k Ž 4K rok to Vout W & 20k lou

OX oßiß2 = 100, VA = 00 a) Find the overall Voltage gaina out = 3909 W b) find the current gain Contain, Rin and Rout νς

€ Repeat exercise a, if the source vs has an internal resistance of ik and ro=50kn for each transistor.

also the given answers might not be correct so don't mind them
Answer #1

Given .dc Equivalents of both Hansistors are : 10: B=B2=100 3 ukr 85k1 3 3 am 80K2 I . 15khi į o.sk vth Rih . 15kr Th= 10x V-0.04273 ofte Bs 100 Im = 0.04864 Boga - 100161 04864 87 = 20055 kn 004273 8= 2.34 kr. Small signal equivalent small signal eI I = 0-14625& ib de Iole in +0.1611 tarb Ioi = {bto 501 = 1-1611ib Iin=86 +0.146asib In = 1.14695ib To = -10016. Tapt = -100K Vol pripovo a) - rooing & Laco Sukr F. 12.75 Kn kn V2 R2 12.75k. V2 = Voix Via Vs* 16KR . 16 Kutiper V1 = 0-4911V Ri= 1K1 1Rin= o. qul Kn+2.34k2 Rin = 3.2811kn Raut = 50km/lake Rout. = 32703k.